# Uplifting Lower-Income Data: Strategies for Socioeconomic Perspective Shifts in Vision-Language Models

### v2

@@ 4 Research Questions
Finally, an interesting finding is that for higher-income images, i.e., up-mid and rich, all income suffixes, including rich and neutral suffixes, lead to drops in Recall. This means that the default English prompts obtain the best results for higher-income images. This could be due to the high representation of higher-income images in AI models and datasets as the “standard representation”. Therefore, including additional information, such as socioeconomic status, causes the model to favor the retrieval of lower-income images over higher-income images. This can be seen in the results that achieve Recall improvements on lower-income images while reducing Recall for higher-income images. Such effects could result from these prompts shifting the VL’s perspective away from its understanding of the default representation of a topic.
@@ 5 Lessons Learned
Current multilingual VL models do not contribute significant improvements to diversity and representation.
Our results from Section 4 demonstrate that English prompts obtain better performance on lower (and higher) income data compared to prompts translated to a non-English that is widely spoken in the region where the data was collected from. Since the quality of translations, quantity of training data available for these languages, and consequently, the performance of AI models in these languages is lower than that of English, these findings are not very surprising. We can look forward to better non-English language performance as multilingual VL models improve.
The addition of extra information to prompts shifts the focus of model inferences from higher-income images.
We find that adding geographical and socioeconomic attributes (including neutral attributes) to prompts leads to an increased model preference for lower-income images over higher-income images as demonstrated in Section 4.
@@ 6 Conclusion
In this paper, we addressed the uneven performance of VL models across different countries and different income levels. We explored three attribute-integrated prompting strategies: (1) translation of text prompts to native non-English languages, (2) addition of geographic information, and (3) addition of socioeconomic attributes. We found that integrating geographical and socioeconomic information into prompts enhances the performance on data from lower-income households and retrieves more diverse label representations. Furthermore, we identified and highlighted the contexts where the proposed prompting techniques work best and shared our insights to improve representation in vision-language models and datasets. Our code can be used to evaluate the performance of other VL models and datasets and is publicly available at Analysis for Uplifting lower-income data.
@@
Our study is constrained by the reach of the Dollar Street dataset and the number of contributions obtained from each region. therefore we do not account for data from other regions not included in the dataset.
@@
Through this work, we aim to contribute toward improving diversity in AI models and even out the disparate impact of these models on the public, especially on underrepresented groups. The strategies discussed in our work can be used to prioritize the retrieval of lower-income images for balancing skewed data representation or domain-specific applications in AI. However, we do not encourage the use of these strategies to promote over-representation or inclusion of one group over the other in contexts that affect all members of the general public.
Our decision to use the NLLB-SigLIP model exemplifies our commitment to using inclusive models that benefit as many people as possible, especially underrepresented groups. While researching technologically advanced communities is easier and less resource-intensive, we stress the importance of making AI design decisions that do not exclude communities with limited access to technology.
@@ Acknowledgements
We are grateful to the Language and Information Technologies (LIT) lab members at the University of Michigan for the insightful discussions and feedback during the early stages of the project. This project was partially funded by a grant from the Department of State (#STC10023GR0014). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the Department of State.
Appendix A Example Appendix
@@ A.2 Lists of countries and classifications
A.3 Recall from all Prompting techniques across all countries
A.4 Non-English Languages and Effect on Recall
@@
A.5 How do different country suffixes prompts affect the Recall of images from different income groups (expanded)?
@@
A.6 Does adding income suffixes improve performance for lower-income images?
A.7 The best-performing country suffixes for higher income data from a continent are from the same continent
| Images by Continent | World Bank Income Classification | Country Suffixes | |||
| African | American | Asian | European | ||
| African images | Poor countries | 42.3 (-1.4) | 40.3 (-3.4) | 36.0 (-7.7) | 40.3 (-3.4) |
| Low-mid countries | 41.8 (-8.8) | 39.7 (-10.9) | 37.0 (-13.6) | 40.7 (-9.9) | |
| Up-mid countries | 33.3 (-5.9) | 32.4 (-6.8) | 25.7 (-13.5) | 33.4 (-5.8) | |
| Average | 39.1 (-5.4) | 37.5 (-7.0) | 32.9 (-11.6) | 38.1 (-6.4) | |
| American images | Up-mid countries | 24.3 (-20.8) | 37.6 (-7.5) | 26.9 (-18.3) | 37.2 (-8.0) |
| Rich countries | 28.7 (-33.4) | 45.7 (-16.4) | 32.5 (-29.6) | 44.7 (-17.4) | |
| Average | 26.5 (-27.1) | 41.7 (-12.0) | 29.7 (-24.0) | 41.0 (-12.7) | |
| Asian images | Low-mid countries | 33.8 (-17.2) | 38.4 (-12.6) | 40.4 (-10.5) | 43.4 (-7.6) |
| Up-mid countries | 30.9 (-21.0) | 40.5 (-11.4) | 41.2 (-10.7) | 46.5 (-5.4) | |
| Rich countries | 28.2 (-19.9) | 36.1 (-12.0) | 36.5 (-11.6) | 40.1 (-8.0) | |
| Average | 31.0 (-19.4) | 38.3 (-12.0) | 39.4 (-10.9) | 43.3 (-7.0) | |
| European images | Low-mid countries | 22.3 (-23.9) | 33.3 (-13.1) | 26.6 (-19.6) | 42.2 (-0.4) |
| Up-mid countries | 18.9 (-25.0) | 29.2 (-14.7) | 24.5 (-19.4) | 40.7 (-3.2) | |
| Rich countries | 19.0 (-20.8) | 27.8 (-12.0) | 21.5 (-18.3) | 37.6 (-2.2) | |
| Average | 20.1 (-23.2) | 30.0 (-13.3) | 24.2 (-19.1) | 40.2 (-3.1) | |
This is a section in the appendix.

### v3

@@ 4 Research Questions
An interesting finding is that all income suffixes, including rich and neutral, result in decreased Recall for higher-income images (i.e., up-mid and rich). This suggests that default English prompts yield the best results for higher-income images, likely due to their high representation in AI models and datasets as the "standard representation." Consequently, the inclusion of socioeconomic status information may lead the model to prioritize lower-income images over higher-income ones. This phenomenon is evident in the results, which show Recall improvements for lower-income images while diminishing Recall for higher-income images, potentially indicating a shift in the model’s perspective away from its default understanding of the topic.
| Model | Suffix Prompts | ||||
| English | Country | Poor | Rich | Neutral | |
| NLLB SigLIP | 30.3 | 41.6 | 32.1 | 30.2 | 29.5 |
| Sentence Transformers MCLIP | 22.2 | 44.9 | 28.6 | 26.7 | 25.2 |
| Open AI CLIP ViT 32/B | 25.4 | 45.0 | 30.0 | 28.1 | 28.3 |
| Model | English | Native Translated |
| NLLB SigLIP | 31.3 | 28.5 |
| Sentence Transformers MCLIP | 24.9 | 18.3 |
4.4 Results Significance and Generalizability
We conducted the Wilcoxon Signed Rank Woolson (2005) test (p-value < 0.05) to assess the statistical significance of our findings. The results indicated that the differences between the default English prompt results and each prompt intervention were statistically significant, except for the ’rich’ and ’neutral’ income suffix prompts (more details in Appendix Table 14).
Although our primary focus is on the NLLB-CLIP-SigLIP results, we confirmed that these findings are consistent across the two other LMM models we tested (Open AI’s CLIP ViT B/32 Radford et al. (2021) and Sentence Transformers clip-ViT-B-32-multilingual-v1 Reimers and Gurevych (2019)). A summary of results from these additional models is included in Table 7 and Table 8.
@@ 5 Lessons Learned
We highlight key insights learned from our findings and present them below.
Current multilingual LMM models do not significantly improve diversity and representation.
Our results from Section 4.1 demonstrate that English prompts perform better on lower (and higher) income images than prompts translated to a non-English language widely spoken in the region where the data was collected. Since the quality of translations, quantity of training data available for these languages, and consequently, the performance of AI models in these languages is lower than that of English, these findings are not very surprising. We can look forward to better non-English language performance as multilingual LMM models improve.
Location and socio-economic attributes improve retrieval performance for lower-income images.
We find that adding geographical and socioeconomic attributes (including rich and neutral attributes) to prompts leads to an increased model preference for lower-income images over higher-income images, as demonstrated in Section 4.2. Images from poor households typically suffer the most from underrepresentation as they differ the most from the type of images available on the internet Rosling et al. (2019). Since LMM models have learned representations from high-income images as the standard, then adding more information to the prompt (such as country suffixes like ’Malawi’, income suffixes describing poverty or wealth, or neutral suffixes like ’a place’) shifts the perspective of the model to retrieve images that are more diverse and less contained to the learned ’standard’.
@@ 6 Conclusion
In this paper, we addressed the uneven performance of LMM models across different countries and income levels. We explored three attribute-integrated prompting strategies: (1) translation of text prompts to native non-English languages, (2) addition of geographic information, and (3) addition of socioeconomic attributes. We found that integrating geographical and socioeconomic information into prompts enhances LMM model performance on images from lower-income households and retrieves more diverse label representations. Furthermore, we identified and highlighted the contexts where the proposed prompting techniques work best and shared our insights to improve representation in LMM models and datasets. Our code can be used to evaluate the performance of other LMM models and datasets and is publicly available at Analysis for Uplifting lower-income data.
@@
Our study is constrained by the reach of the Dollar Street dataset and the number of contributions obtained from each region. Therefore, we do not account for data from other regions that are not included in the dataset.
@@
Through this work, we aim to contribute toward improving diversity in AI models and even out the disparate impact of these models on the public, especially on underrepresented groups. The strategies discussed in our work can be used to prioritize the retrieval of lower-income images for balancing skewed data representation or domain-specific applications in AI. However, we do not encourage the use of these strategies to promote over-representation or the inclusion of one group over another in contexts that affect all members of the general public.
Our decision to use the NLLB-SigLIP model exemplifies our commitment to inclusive models that benefit as many people as possible, especially underrepresented groups. While researching technologically advanced communities is easier and less resource-intensive, we stress the importance of making AI design decisions that do not exclude communities with limited access to technology.
@@ Acknowledgements
We are grateful to the Language and Information Technologies (LIT) lab members at the University of Michigan for their insightful discussions and feedback during the project’s early stages. This project was partially funded by a grant from the Department of State (#STC10023GR0014). Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the Department of State.
Appendix A Appendix
@@ A.1 Non-English Languages
| Language | ISO | chrf++ |
| Arabic | arb_Arab | 51.4 |
| Bengali | ben_Beng | 46.2 |
| Burmese | mya_Mymr | 29.3 |
| Chinese | zho_Hans | 19.6 |
| Creole | hat_Latn | 50.2 |
| Czech | ces_Latn | 52.7 |
| Danish | dan_Latn | 63.2 |
| Dutch | nld_Latn | 53.1 |
| Ewe | ewe_Latn | 35.6 |
| Farsi_Persian | pes_Arab | 47.4 |
| French | fra_Latn | 67.0 |
| German | deu_Latn | 59.4 |
| Hausa | hau_Latn | 49.0 |
| Hindi | hin_Latn | 54.2 |
| Indonesian | ind_Latn | 66.6 |
| Italian | ita_Latn | 54.6 |
| Khmer | khm_Khmr | 31.2 |
| Kinyarwanda | kin_Latn | 44.0 |
| Korean | kor_Hang | 32.1 |
| Kyrgyz | kir_Cyrl | 42.6 |
| Mongolian | khk_Cyrl | 37.3 |
| Nepali | npi_Deva | 49.0 |
| Oromo | gaz_Latn | 31.6 |
| Portuguese | por_Latn | 67.4 |
| Romanian | ron_Latn | 58.2 |
| Russian | rus_Cyrl | 52.5 |
| Serbian | srp_Cyrl | 53.3 |
| Shona | sna_Latn | 42.9 |
| Sinhala | sin_Sinh | 42.4 |
| Somali | som_Latn | 41.5 |
| Spanish | spa_Latn | 52.6 |
| Swahili | swh_Latn | 58.0 |
| Swedish | swe_Latn | 62.7 |
| Tagalog | tgl_Latn | 56.4 |
| Thai | tha_Thai | 36.0 |
| Turkish | tur_Latn | 52.9 |
| Ukrainian | ukr_Cyrl | 50.5 |
| Urdu | urd_Arab | 46.6 |
| Vietnamese | vie_Latn | 56.4 |
| Zulu | zul_Latn | 51.0 |
| Prompt | P-value | Sig. or not |
| English & Native translated | 8.64e-09 | yes |
| English & Country suffix | 7.27e-08 | yes |
| English & Poor Income Suffix | 0.02 | yes |
| English & Rich Income Suffix | 0.603 | no |
| English & Neutral Income Suffix | 0.563 | no |